OpenStudy (anonymous):

Simplify log3(1/27)^4 Would this simplify down to 4 log3^1 - 4log3^3

OpenStudy (tkhunny):

?? The "3" is a Base, not an argument. \(1/27 = 3^{-3}\)

OpenStudy (anonymous):

4 log3 1 - 4 log3 3^(-3)

OpenStudy (tkhunny):

You missed the first hint. \(\log_{3}(a)\) has only a little to do with \(\log(3)\) \(\log_{3}(1/27)^{4} = 4\cdot\log_{3}(1/27) = 4\cdot\log_{3}(3^{-3}) = -12\cdot\log_{3}(3)\) Can you finish? Only one more thing to do.

OpenStudy (anonymous):

What do I need to do to finish? 4 log3 1 - (-12 log3 (3)?

hartnn (hartnn):

use the fact that log 1 in any base = 0 and \(\log_aa=1\) so, \(\log_33=1\)

hartnn (hartnn):

how did you get 2 negatives ?

hartnn (hartnn):

27 is \(3^3\) and not \(3^{-3}\)

OpenStudy (anonymous):

Ok so am I looking at 4 log3 1-12 or have I gotten my self so confused I' way off ?

hartnn (hartnn):

so you used \(\log_33=1\) and thats correct! now also use the fact that \(\log_31=0\)

OpenStudy (anonymous):

So my answer is 0-12 ?

hartnn (hartnn):

which is -12 and yes :)

OpenStudy (anonymous):

Finally!!!! Thank you ever so much!

hartnn (hartnn):

you're most welcome ^_^